# Culturally Aware and Adapted NLP: A Taxonomy and a Survey of the State of the Art

### v1

@@ 5 Creating Culturally Adapted Models
for tree= forked edges, grow’=0, draw, rounded corners, node options=align=center,, text width=2cm, s sep=4pt, calign=edge midpoint, l sep=7.5pt, , [, phantom [Model §5, finetuning_pt [Data Aug., finetuning [Li and Zhang (2023), finetuning_work] ] [Cont. Pre-training; Auxilary Losses, finetuning [Hofmann et al. (2024); Li et al. (2024), finetuning_work] ] [Information Integration, finetuning [Yao et al. (2023); Friedrich et al. (2023), finetuning_work] ] [Parameter-Efficient Adaptations, finetuning [Xiao et al. (2023); Liu et al. (2023d), finetuning_work] ] ] ]
@@
5.1 Data Augmentation
It can be challenging to acquire large corpora for supervised training in cultural adaptation. Data augmentation has been supporting this process and increasing the robustness of models. Li and Zhang (2023) present a data augmentation technique (CultureMixup) for multilingual multicultural VL reasoning tasks. CultureMixup creates code-mixed data with concepts in English replaced through cultural concept mapping for supervised training. The cultural concept sets (for concept mapping) are constructed by querying hyponyms, synonyms, and hypernyms in the ConceptNet Speer et al. (2017) and WordNet Miller (1995). However, the optimal resource will depend on the cultural element in question (§3.1). For example, a cultural knowledge base might be better for norms adaptations.
5.2 Continual Pre-training, Auxiliary Losses
Continual pre-training (CPT, including instruction tuning), intermediate task training, and utilizing multi-task or auxiliary losses have also been used for cultural adaptation. CPT involves tuning a pre-trained LM with an additional unlabeled domain or language corpus before downstream task fine-tuning. It can improve downstream task performance via full-parameter training Xu et al. (2019); Han and Eisenstein (2019); Gururangan et al. (2020) or by training a small set of additional parameters while keeping the original model frozen Pfeiffer et al. (2020); Wang et al. (2021); Ke et al. (2022).
Recently, Hofmann et al. (2024) show that when combined with a geo-location prediction loss, CPT can help to increase the awareness of dialectal variations of pre-trained LMs. Wang et al. (2024) show that mining instructions containing cultural knowledge and performing instruction tuning can improve models’ ability in cultural knowledge reasoning. In VL settings, Bhatia and Shwartz (2023) use a cultural common sense knowledge graph from  Nguyen et al. (2023) for CPT, developing a geo-diverse LM for downstream commonsense reasoning tasks. CultureLLM Li et al. (2024) reformulates the World Value Survey as an intermediate training task. It shows that LLMs can enhance their downstream performance in multicultural hate speech classification where diverse value perceptions across cultures are required. However, adapting LLMs to culture may result in catastrophic forgetting (McCloskey and Cohen 1989, or termed “alignment tax” due to RLHF tuning, Ouyang et al. 2022; Askell et al. 2021), potentially worsening their performance on general tasks. Similarly, general continual fine-tuning may shift existing cultural values Choenni et al. (2024) in pretrained models. These warrant further investigation.
5.3 Other Forms of Information Integration
Missing cultural context, knowledge or demographic information have also been integrated into models in other ways:
As representations. Cao et al. (2024a) introduce the cuDialog dataset and models for multi-turn dialogue classification and prediction through a cultural lens. Using an encoder-decoder Transformer (such as mT5,  Xue et al. 2021), the cultural dimensions vectors are concatenated to the hidden states at each layer. The cultural context vectors were obtained through a regression task over the encoder hidden states with dialogue inputs to predict the corresponding cultural scores (based on Hofstede Culture Dimensions, Hofstede 1984).
As part of the prompt. AlKhamissi et al. (2024) enhance LLMs’ cultural alignment by integrating anthropological reasoning with demographic-based prompts. In a game setting with multicultural players,  Shaikh et al. (2023) encode sociocultural attributes into the input to enrich the model’s ability for game plays. Similarly, Yao et al. (2023) propose strategies to enhance the cultural translation ability of LLMs with reference-explanation of cultural entities. LLMs effectively leverage indirect descriptions obtained from external data sources or generated by LLMs as a prior step.
Friedrich et al. (2023) propose a retrieval-based method for moral reasoning tasks. Moral contexts specific to a culture (or user) are stored in a retrieval engine. When querying with moral questions, relevant contexts are retrieved and prepend to the query as input to LLMs. LLMs can change their default answer based on the retrieved context. Using a retrieval-augmented system can be a promising future direction for adapting rich and evolving cultural information.
5.4 Parameter-Efficient Adaptations
As LMs grow larger, parameter-efficient fine-tuning methods (i.e., PEFT, by fine-tuning a small number of parameters, such as the bottle-neck adapters, Houlsby et al. 2019; LoRA, Hu et al. 2022 etc.) become increasingly important for task adaptations. Given their success in cross-lingual transfer learning (Pfeiffer et al., 2020; Ansell et al., 2021; Liu et al., 2023a, c; Üstün et al., 2020, among others), PEFT can be a natural choice for cultural adaptation of e.g., dialects.
Recently, HyperLoRA Xiao et al. (2023) uses the Hypernetwork (Ha et al., 2017, a neural network for generating parameters) to generate LoRA adapters based on dialectal features. DADA Liu et al. (2023d) proposes to train a pool of dialectal linguistic feature adapters and dynamically compose the adapters for dialectal tasks. Being task agnostic, PEFT methods could prove important for cultural adaptations beyond dialects.
5.5 Outlook: Reinforcement Learning from Human Feedback (RLHF), Direct Preference Optimization (DPO)
The success of LLMs has popularized RLHF Christiano et al. (2017); Bai et al. (2022); Ouyang et al. (2022) and DPO Rafailov et al. (2023); Ivison et al. (2023) methods. RLHF fine-tunes LMs with feedback by fitting a reward model with human preferences, and then training a reinforcement learning-based policy to maximize the learned reward. DPO avoids RL training by using a simpler supervised learning objective for an implicit reward model.
Recent work shows that RLHF can enhance the performance of multilingual instruction tuning for LLMs Lai et al. (2023), while DPO can improve the multilingual reasoning abilities She et al. (2024) of LLMs. The use of RLHF or DPO for multilingual multicultural adaptation is still limited, but these examples suggest that the direction could be promising.
6 Further Discussions
As we have seen, “culture” emerges as a highly active area of NLP research. Yet significant work remains to be done on both resources and methods for various elements of culture.
An area that requires attention is the overall process of researching culturally aware NLP. First, it is important to consider whether and how the target communities could best benefit from cultural technologies at this point in time. For example, many dialects are primarily oral, and speech-to-speech or speech-to-text translations may be preferable over text-based applications Blaschke et al. (2024). Or, if the culture in question does not have LLMs with basic conversational fluency yet, it may be more important to address this problem first, before embarking on improving pragmatic alignments of LLMs for the culture. Getting the process right requires consultation with target communities Bird (2020); Liu et al. (2022b); Mager et al. (2023). Ethical data collection practices are also critical and technology ownership must be considered. We refer the readers to work such as Bird (2020); Smith (2021); Cooper et al. (2024) for further details of this.
Another area to consider is the integration of insights and practices from fields beyond NLP. Incorporating cultural factors and performing cultural adaptations has a long history in application areas, such as for video games O’hagan and Mangiron (2013), movies Pettit (2009), online learning systems Blanchard et al. (2005), and clinical psychology intervention Bernal et al. (1995); Barrera Jr and Castro (2006). Existing practices should be used as a foundation for adapting NLP applications to meet the needs of diverse cultural contexts.
7 Summary and Future Research Directions
Culturally aware and adapted NLP has recently emerged as a highly active research area. Significant progress has been made in the development of resources for capturing various elements of culture, but the development of NLP methods is still in its infancy. We will now summarize the main research gaps identified in this survey with respect to the categories of our new taxonomy (§2 for details):
Cultural: Multilingual data resources covering a diverse set of concepts (e.g., aesthetics, spatial relation) in both unimodal and multimodal (§3.1.1) for generation tasks are needed. Moreover, most recent developments in norms & morals are predominantly in English and reflect a monocultural perspective, which indicates the need for further development of multilingual and multicultural resources. Additionally, there is a significant gap in datasets that focus on different types of value perceptions (such as emotion and irony, §3.1.3), stylistic variations (§3.1.5), and artifacts (§3.1.6) across various cultural groups both in different languages and within languages.
Sociocultural: Resources that explicitly consider sociocultural elements of culture (§3.2) are still lacking. For example, collecting speaker relationships in dialogue datasets or distinguishing age groups in social norms datasets. These are needed to address the intricate relationship between culture and people in NLP.
Training Data and “CultureGLUE”: Most existing resources focus on evaluation, but there is a pressing need for training data. In addition, there is no unified cultural equivalent of the GLUE benchmark that encompasses all elements of culture across a wide range of cultures. Creating a multicultural “CultureGLUE” may be challenging at the moment, a good first step is to focus on individual cultures with diverse tasks and element coverage.
Modelling. While modelling methods for culture are generally under-explored, continual pretraining (§5.2) and prompting (§5.3) have received marginally more attention than other approaches. Research areas needing further exploration include:
PEFT-based Transfer Learning: Exploration of PEFT-based transfer learning techniques beyond dialects is limited (§5.4). Given their success in other NLP areas, these techniques warrant further investigation into other elements of culture, such as values or norms & morals.
RLHF, DPO and Other LLMs Specialties: Leveraging the success of LLMs, RLHF and DPO as a new promising avenue for cultural adaptation (§5.5). It would also be interesting to explore other unique abilities of LLMs (such as tool-use, role-playing etc.) for enabling culturally aware NLP.
Evolving Culture: Culture (slowly) evolves Boyd and Richerson (1988); Whiten et al. (2011), yet there have been little discussions on how to model and adapt to evolving culture. One potential approach is the use of retrieval-augmented systems to integrate evolving information (§5.3), which ensures models’ relevance to cultural shifts over time.
Overall - “Surface” versus “Deep” culturally adapted NLP. Resnicow et al. (1999) devise cultural adaptations for public health research into surface and deep adaptations, where the former considers familiar languages and concepts to the target groups, and the latter considers social and historical factors that influence the behaviours of the target groups. In NLP, surface adaptations might include using the same language as a culture and recognizing explicit cultural differences (e.g., asking LLMs “what is the meaning of …”). In contrast, deep adaptations might enable a model to “behave” (e.g., make decisions, pragmatically comply etc.) like a member of a culture without explicit inquisition (see Figure 4 in the Appendix for an illustration).
Only a few current works focus on adapting the behavioural aspect of models (which is becoming increasingly important with LLMs), and there has been no work to date measuring the depth and progress of cultural adaptations or when a model is fully culturally aware and culturally competent. Further research could explore these areas.
8 Conclusions
This paper proposes a new extensive taxonomy of culture that expands on earlier works in NLP and is grounded in well-established anthropology literature. The taxonomy provides a systematic framework for understanding and tracking progress in the emerging area of culturally aware and adapted NLP. We survey existing resources and methods in this area according to the taxonomy classes, identifying areas of strength as well as areas where research remains to be done. Our paper summarizes the state of the art and provides ideas for future development in this exciting and important area of research.
9 Limitations
In this work, we may only discuss research with respect to a single element of culture. Research related to culture can be multifaceted, considering multiple elements of culture simultaneously. Additionally, our survey emphasizes resources and adaptation methods due to the scope, without extensive coverage of evaluation and probing techniques.
Our taxonomy provides a systematic framework for understanding and tracking progress in the emerging area of culturally aware and adapted NLP. However, our taxonomy is not without its limitations. Future research could refine the taxonomy in areas like values or communicative goals, create subcategories and enhancing our understanding of how elements of culture interact (e.g., how changing values impact social norms over time).
Acknowledgements
This work was funded by the German Federal Ministry of Education and Research (BMBF) under the promotional reference 13N15897 (MISRIK). This work has also been supported by the UK Research and Innovation (UKRI) Frontier Research Grant EP/Y031350/1 EQUATE (the UK government’s funding guarantee for ERC Advanced Grants) awarded to Anna Korhonen at the University of Cambridge. Chen Cecilia Liu is supported by the Konrad Zuse School of Excellence in Learning and Intelligent Systems (ELIZA) through the DAAD program Konrad Zuse Schools of Excellence in Artificial Intelligence, sponsored by the Federal Ministry of Education and Research.
We thank Anjali Kantharuban, Shun Shao, and Anne Lauscher for an early discussion of different aspects of this work. We thank Ji-Ung Lee for feedback on a draft of this paper.
Appendix A Method
We prioritize main and findings papers from ACL, EMNLP, EACL, NAACL and TACL published in 2022, 2023 and early 2024. The search terms used are: “culture”, “cultural”, “geo-diverse”, “socio”, “social”, “moral”, “bias”, “norms” and we prioritize works that focus on datasets and modelling methods for cultural adaptation. Additional relevant papers from other sources are also discussed in this work to a lesser extent and coverage. Evaluation method and probing method papers are discussed to a much lesser extent due to the scope of our paper. We also refer readers to recent surveys like Adilazuarda et al. (2024) for focused discussions.

### v2

@@ Appendix A Method
for tree= forked edges, grow’=0, draw, rounded corners, node options=align=center,, text width=2cm, s sep=4pt, calign=child edge, calign child=(n_children()+1)/2, l sep=7.5pt, , [, phantom [Data §4, data_pt [Manual, data[New, data_sub[ Liu et al. (2021)
Thapliyal et al. (2022)
Felkner et al. (2023)
Koto et al. (2024b)
Lee et al. (2024b), data_work]] data[Adapt, data_sub[ Majewska et al. (2023)
Koto et al. (2023)
Hu et al. (2023), data_work] ] ] [Automatic: Models and Pipelines, data[New, data_sub[Nguyen et al. (2023)
Fung et al. (2023)
Fung et al. (2024), data_work]] data[Adapt, data_sub[Putri et al. (2024), data_work]] ] [Semi-Automatic: Model-in-the-loop; Structured Resources, data[New, data_sub[Ziems et al. (2023a)
CH-Wang et al. (2023)
Liu et al. (2024a)
Bhutani et al. (2024) , data_work]] data[Adapt, data_sub[ Held et al. (2023)
Xiao et al. (2023) , data_work]] ] ] ]
Figure 4: Categorization of the methods for resource acquisitions with representative examples.
{forest}
for tree= forked edges, grow’=0, draw, rounded corners, node options=align=center,, text width=2cm, s sep=4pt, calign=edge midpoint, l sep=7.5pt, , [, phantom [Modelling §5, finetuning_pt [In-Context, finetuning_pt [Sociodemographics Prompt; Cultural Role-playing, finetuning[ Shaikh et al. (2023)
Hwang et al. (2023)
AlKhamissi et al. (2024) , finetuning_work] ] [Cultural Knowledge, finetuning[ Yao et al. (2024a)
Zhou et al. (2024)
, finetuning_work] ] [Reasoning Prompt;
Constitution, finetuning[ Hayati et al. (2024)
AlKhamissi et al. (2024)
Vijjini et al. (2024)
, finetuning_work] ] [Ensemble of LLMs, finetuning[ Feng et al. (2024)
, finetuning_work] ] [Retrieval-Augmented, finetuning[ Friedrich et al. (2023)
Conia et al. (2024)
Hu et al. (2024) , finetuning_work] ] ] [In-Weight, finetuning_pt [Data Augmentation, finetuning [Li and Zhang (2023), finetuning_work] ] [Continual Pre-training; Auxilary Losses, finetuning [ Bhatia and Shwartz (2023)
Zhou et al. (2023b)
Hofmann et al. (2024a)
Spliethöver et al. (2024)
Wang et al. (2024a)
Cahyawijaya et al. (2024)
Kim et al. (2024)
Alyafeai et al. (2024)
, finetuning_work] ] [Information Integration (with vectors), finetuning [ Cao et al. (2024a)
White et al. (2024) , finetuning_work] ] [Parameter-Efficient Adaptations, finetuning [Xiao et al. (2023)
Liu et al. (2023b), finetuning_work] ] [Feedback Learning, finetuning [Aakanksha et al. (2024) , finetuning_work] ] ] ] ]